# State-witness contraction

### v1

@@ Abstract
We present a method to construct entanglement witnesses for arbitrarily large multipartite systems, by tensoring and partial tracing existing states and witnesses. As a proof of principle we show that, using little shared quantum resources, the method allows to reuse witnesses unable to detect states with local positive partial transpositions into new ones able to do so. Moreover, we show that this technique allows to tailor both linear and nonlinear witnesses to specific states using semidefinite programming with comparatively low-dimensional variables. As an example, existing trace polynomial witnesses are significantly improved while preserving their symmetries and implementability with randomized measurements. Besides detecting entanglement, the method is shown to detect $k$-copy distillability. A recipe for the single-copy case is shown to be effective for generic and Werner states.
I Introduction
@@
In the last decades, a large improvement has been done in the control over mutlipartite and high-dimensional quantum systems [1, 2]. This paves a road towards realistic quantum software [3], quantum networks [4, 5] and quantum devices where entanglement constitutes a major nonclassical resource. As a result, detecting the presence of entanglement in multipartite quantum systems becomes crucial [3, 6], but it is at the same time a theoretical and experimental challenge due to the growth of the state space with the local dimension [7, 8].
@@
For this purpose, a vastly used technique is the evaluation of entanglement witnesses, namely observables whose negative expectation value detects the presence of entanglement [9, 10, 11]. However, important challenges remain: for some states, finding a witness detecting them requires unaccessible computational resources [12, 13]; finding new entanglement witnesses for multipartite systems of interest is challenging [14, 15, 16]; and adjusting a witness to a state cannot be done systematically in general [7]. A promising tool detecting larger sets of states [17] are nonlinear witnesses acting on multiple copies [18, 19]. While these overcome some limitations of linear witnesses [17, 20], problems like tailoring nonlinear witnesses to states are even more challenging than in the linear case.
Here we provide a new method to derive new witnesses using existing states and witnesses (Prop. 1 and Fig. 1). We show that existing witnesses can be lifted to larger ones with further detection capabilities (Observation 2). Naturally, the construction given in this work allows to partially adjust specific witnesses to desired states efficiently – see Eq. (4). As a very special case, the method allows to construct multicopy witnesses, which are shown to detect $k$-copy distillability under some conditions. We compare this method with trace polynomial multicopy witnesses we previously introduced in [20], over which significant advantage is shown with little computational effort.
II State-witness contraction
@@
The following result provides a systematic recipe to construct new entanglement witnesses from existing states and witnesses (depicted in Fig. 1). A deeper analysis in terms of positive maps in the special case of trace polynomials is given in [21].
Proposition 1 (State-witness contraction).
@@
Let $\{R^{(j)}\}_{j=1}^{n}$ be $k_{j}$-partite states or witnesses. Let $\tau$ be a $n$-partite state or witness. Then, the operator
[…]
@@
has a nonnegative expectation value, $\operatorname{tr}(\mathcal{W}_{\tau}\varrho)\geq 0$, on separable $(k_{1}+...+k_{n}-n)$-partite states $\varrho$.
@@
This can be seen from the duality between the block-positive cone and the cone of separable states [22] as follows.
@@
Using the coordinate-free definition of the partial trace, for any state $\sigma^{(j)}_{k}$ and product state $\varrho^{(j)}=\varrho_{1}^{(j)}\otimes...\otimes\varrho_{k-1}^{(j)}$ we have $0\leq\operatorname{tr}(R^{(j)}\cdot\varrho^{(j)}\otimes\sigma^{(j)}_{k})=\operatorname{tr}(\theta^{(j)}\otimes\sigma^{(j)}_{k})$, where we define $\theta^{(j)}:=\operatorname{tr}_{1...k-1}(R_{j}\cdot\varrho_{1}^{(j)}\otimes...\otimes\varrho_{k-1}^{(j)}\otimes\mathds{1})$. Equivalently, $\theta^{(j)}$ is positive semidefinite for all $1\leq j\leq n$. Therefore, for any $n$-partite state or witness $\tau$, we have the inequality $0\leq\operatorname{tr}(\theta^{(1)}\otimes...\otimes\theta^{(n)}\cdot\tau)=\operatorname{tr}(\mathcal{W}_{\tau}\varrho)$. Here the equality is obtained by plugging in the definition (1) of $\mathcal{W}_{\tau}$, and we assume $\varrho:=\varrho^{(1)}\otimes...\otimes\varrho^{(n)}$ where each $\varrho^{(j)}$ is a product $(k_{j}-1)$-partite state. By convexity, $\operatorname{tr}(\mathcal{W}_{\tau}\varrho)\geq 0$ also holds for any separable $(k_{1}+...+k_{n}-n)$-partite state $\varrho$.
@@
Figure 1: State-witness contraction yielding a four-partite witness from two three-partite operators. Here we take the tensor product of a witness $W$ and a positive operator $M$. Then, instead of obtaining a multicopy or trace polynomial witness as in [20, 23], the first (or last, equivalently) subsystems $A_{1}$ and $B_{1}$ are contracted (multiplied and partial traced) with a state or witness $\tau_{A_{1}B_{1}}$ to be chosen – see Eqs. (1) and (3). This allows to lift $W$ to a larger witness $\mathcal{W}_{\tau}$ with further detection capabilities (Observation 2), which moreover can be partially optimized over $\tau$ to detect specific four-partite states – see Eq. (4).
III Locally bound entanglement
@@
As a case of interest, we will employ this method to reuse existing highly symmetric witnesses given in [23], which belong to a larger class of methods involving permutations in the Hilbert space [24, 25]. While these are desired due to their symmetries and implementation, their use in detecting states with local positive partial transpositions is a current challenge [26, 14, 23, 27, 28, 29, 30, 31]. In particular, consider the witness
@@
constructed from the so-called Young projectors $P_{2,1}=\eta_{d}\big{(}2(\operatorname{id})-(123)-(132)\big{)}/6$ and $P_{1^{3}}=\eta_{d}\big{(}(\operatorname{id})-(12)-(13)-(23)+(123)+(132)\big{)}/6$, where $\eta_{d}(\pi)\mathinner{|v_{1}\rangle}\mathinner{|v_{2}\rangle}\mathinner{|v_{3}\rangle}=\mathinner{|v_{\pi^{-1}(1)}\rangle}\mathinner{|v_{\pi^{-1}(2)}\rangle}\mathinner{|v_{\pi^{-1}(3)}\rangle}\in{\mathds{C}^{d}}^{\otimes 3}$ represents the permutation $\pi\in S_{3}$ onto ${(\mathds{C}^{d})^{\otimes 3}}$. This witness can be derived from immanant inequalities [32, 33] using the techniques in [23, 26, 20]. Linear optimization shows that it can detect three-partite states with zero, one or two positive partial transpositions, but not three (see Table II in [21]). Using Proposition 1 with an ancillary Bell state $\mathinner{|\phi^{+}\rangle}=\sum_{i=0}^{2}\mathinner{|ii\rangle}/\sqrt{3}$, it can be recycled to detect four-partite states whose local parties have positive partial transpositions, and thus cannot be distilled with local operations with classical communication if the four parties are spatially separated. This is a type of unlockable bound entanglement [34, 35, 36, 37, 38, 39], which we refer to as locally bound.
@@
can detect four-partite entangled states $\varrho_{A_{1}A_{2}B_{1}B_{2}}$ with $\varrho^{T_{i}}\geq 0$ for $i\in\{A_{1},A_{2},B_{1},B_{2}\}$. In particular, $\min\{\operatorname{tr}(\mathcal{W}_{\tau}):\varrho^{T_{i}}\geq 0\}\approx-0.000276$
Note that this can be achieved using the shared quantum resources provided by the Bell state in an ancillary system $A_{3}B_{3}$. Yet, these can be reduced significantly: using a less entangled state $\mathinner{|\phi_{s}\rangle}=\sqrt{s}\mathinner{|00\rangle}+\sqrt{1-s}\mathinner{|11\rangle}$ embedded into $\mathds{C}^{3}\otimes\mathds{C}^{3}$ with $s=0.9999$, a minimum value $\min_{\varrho^{T_{i}}\geq 0}\langle\mathcal{W}_{\tau}\rangle_{\varrho}\approx-6.051\cdot 10^{-5}$ is obtained using only $-s\log_{2}(s)-(1-s)\log_{2}(1-s)\approx 0.0015$ ebits of entanglement.
IV Witnesses tailored to states
To some degree, witnesses constructed in Proposition 1 can be efficiently adjusted to specific states: consider a fixed set of $n$ $k_{j}$-partite witnesses (or states) $R^{(i)}$ and a target $(k_{1}+...+k_{n}-n)$-partite state $\varrho$ in local dimension $d$ to be detected. Let us choose $\tau$ to be optimized over all possible states and decomposable witnesses [9],
@@ IV Witnesses tailored to states
Here $\mathcal{S}$ denotes any possible subset of the last subsystems $\{k_{1},...,k_{n}\}$. While by Proposition 1 one can choose $\tau$ to be any block-positive operator (any state or witness), the optimization becomes accessible with semidefinite programming if $\tau$ is restricted to be a decomposable witness or a state [40]. Notice that the variable $\tau$ is of size $d^{n}$ [41], even though the target state to be detected is described by a density matrix of size $d^{(k-1)n}$ (assuming $k_{1}=...=k_{n}:=k$ for simplicity).
@@
$\mathcal{W}_{\tau^{*}}=\operatorname{tr}_{A_{2}B_{2}}\Big{(}\eta_{d}(12)_{A_{1}B_{1}}\otimes\phi^{+}_{B_{1}B_{2}}\cdot\varrho_{A_{1}B_{1}}\otimes\tau^{*}_{A_{2}B_{2}}\Big{)}$ (5)
with $\tau^{*}=\text{argmin}_{\tau}\big{\{}\operatorname{tr}(\mathcal{W}_{\tau}\varrho):\tau\geq 0,\operatorname{tr}\tau=1\big{\}}$ detects any state $\varrho$ with nonpositive partial transpose. Similarly, standard witness tailoring techniques using diagonalization [9] are recovered with this example: the duality between positive maps and witnesses [42] is used to detect any state detected by a positive map $\Lambda_{A}\otimes\operatorname{id}_{B}(\varrho_{AB})\not\geq 0$, replacing $\phi^{+}$ in Eq. (5) for $\Lambda^{\dagger}\otimes\operatorname{id}(\eta_{d}(12))$ where $\Lambda^{\dagger}$ satisfies $\operatorname{tr}(\Lambda(X)Y)=\operatorname{tr}(X\Lambda^{\dagger}(Y))$ for all operators $X$ and $Y$.
V Multiple copies
Here we will focus on the case where the state to be detected consists of multiple copies $\varrho^{\otimes k}$ of a target state $\varrho$. We shall demonstrate that the technique introduced in this work allows to significantly improve our previous trace-polynomial multicopy constructions given in [20]. The latter are obtained as a very special case of the constructions given by Proposition 1, by setting the input state to be composed of $k-1$ copies of a $n$-partite target state $\varrho^{\otimes k-1}$. For example, consider a Bell state affected by white noise,
@@ V Multiple copies
which is often used as a worse-case scenario in noisy setups [43]. In [20] we showed that this state can be detected for a range of values of $p$ with the nonlinear witness $\mathcal{W}=(\mathds{1}-6P_{1^{3}})_{A_{123}}\otimes(P_{1^{3}})_{B_{123}}$ acting on three copies of $\varrho$, through $0<\operatorname{tr}(\mathcal{W}\varrho^{\otimes 3})=\operatorname{tr}(\mathcal{W}_{\varrho}\varrho^{\otimes 2})$ where we denote $\mathcal{W}_{\varrho}:=\operatorname{tr}_{A_{3}B_{3}}(\mathcal{W}\cdot\mathds{1}\otimes\varrho_{A_{3}B_{3}})$ for easier comparison to our present work. From Proposition 1 it becomes clear that $\mathcal{W}$ can be better tuned to $\varrho$ through Eq. (4), in this case leading to (Fig. 1)
$\mathcal{W}_{\tau}:=\operatorname{tr}_{A_{3}B_{3}}\Big{(}\mathcal{W}\cdot(\mathds{1}_{A_{1}A_{2}B_{1}B_{2}}\otimes\phi^{+}_{A_{3}B_{3}})\Big{)}\,.$ (7)
@@
Figure 2: Tailoring multicopy witnesses to states. Linear witnesses detect regions of entangled states separated by hyperplanes (red area) through $\langle W\rangle_{\varrho}=\operatorname{tr}(W\varrho)<0$. Multicopy witnesses $\mathcal{W}$ detect regions of entangled states through $\langle\mathcal{W}\rangle_{\varrho^{\otimes k}}<0$ [18, 19], separated by curved hypersurfaces (magenta and blue areas). In particular, $\mathcal{W}$ can be constructed by tensoring linear witnesses [20]. By replacing one of the copies by a variable state or witness $\tau$, the state-witness contraction technique (SWC) allows to systematically detect a larger set of states (green area –see Table 1 for an example) and approach further the set of separable states (SEP).
@@
Observation 3.
Let $\mathcal{W}_{\tau}$ in Eq. (4) be constructed from trace polynomial operators $R^{(j)}$ obtained by immanant inequalities according to [23]. Then, the expression $\operatorname{tr}(\mathcal{W}_{\tau}\varrho^{\otimes k-1})$ is invariant under local unitaries acting on $\varrho$.
@@
Local unitaries act as $U^{\otimes k_{j}-1}$ onto the first $k_{j}-1$ copies of each $j$’th local party. A minimum obtained for $\varrho_{AB}$ by $\tau_{AB}$ is also obtained for $U\otimes V\varrho_{AB}U^{\dagger}\otimes V^{\dagger}$ by $U\otimes V\tau_{AB}U^{\dagger}\otimes V^{\dagger}$ (extension to the multipartite case is straightforward). Thus local unitaries act in practice as $U^{\otimes k_{j}}$ on the $j$’th party. The rest of the argument is the same as in [20], namely $U^{\otimes k_{j}}$ commutes with the representation $\eta_{d}(\pi)$ of permutations $\pi\in S_{k_{j}}$ due to the Schur-Weyl duality. ∎

### v2

# State – witness contraction

@@ Abstract
We construct both nonlinear and linear entanglement witnesses, by tensoring and partial tracing existing states and witnesses. We show that little shared quantum resources allow to employ decomposable witnesses to obtain larger ones detecting locally undistillable states, and find analytic witnesses for multipartite entangled states that are undistillable across all bipartitions. Crucially, we show how to efficiently optimize multicopy witnesses to desired states, which is a current challenge. As an example of both theoretical and experimental interest, existing trace polynomial witnesses are significantly improved while preserving their symmetries and implementability with randomized measurements. Besides detecting entanglement, we also find new witnesses detecting $k$-copy distillability. A recipe for the single-copy case is shown to be effective for generic and Werner states.
@@
Introduction.   In the last decades, a large improvement has been done in manipulating mutlipartite and high-dimensional quantum systems [1, 2]. This paves a road towards realistic quantum software [3], quantum networks [4, 5] and quantum devices where entanglement constitutes a major nonclassical resource. As a result, detecting the presence of entanglement in multipartite quantum systems becomes crucial [3, 6], but it is at the same time a theoretical and experimental challenge due to the growth of the state space with the local dimension [7, 8].
@@
For this purpose, a vastly used technique is the evaluation of entanglement witnesses, namely observables whose negative expectation value detects the presence of entanglement [9, 10, 11]. However, important challenges remain: for some states, finding a witness detecting them requires unaccessible computational resources [12, 13]; finding new entanglement witnesses for multipartite systems of interest is challenging [14, 15, 16]; and adjusting a witness to a state cannot be done systematically in general [7]. A promising tool detecting larger sets of states [17] are nonlinear witnesses acting on multiple copies [18, 19]. While these overcome some limitations of linear witnesses [17, 20], problems like tailoring multicopy witnesses [18, 19, 20] to states are even more challenging than in the linear case.
Here we provide a method to derive both linear and nonlinear entanglement witnesses using existing states and witnesses (Prop. 1 and Fig. 1). We show that little shared entanglement suffices to lift existing witnesses to larger ones with further detection capabilities (Observation 2), and we construct simple witnesses detecting bound multipartite entanglement (Proposition 3). By adjusting the constructed witnesses – see Eq. (6), we are able to optimize to target states a class witnesses acting on multiple copies [18, 19] (Fig. 2 and Table 1). Besides detecting entanglement, the method is shown to be effective at detecting $k$-copy distillability – see Eq. (10).
@@
State-witness contraction.   The following result provides a systematic recipe to construct new entanglement witnesses from existing states and witnesses (depicted in Fig. 1).
Proposition 1.
@@
Let $\{R^{(j)}\}_{j=1}^{n}$ be states or witnesses in $k_{j}$-partite systems $S^{j}$. Let $\tau_{S^{1}_{k_{1}}...S^{n}_{k_{n}}}$ be a $n$-partite state or witness shared between the last parties of all systems $S^{j}$. Then, the operator
$\mathcal{W}_{\tau}:=\operatorname{tr}_{S^{1}_{k_{1}}...S^{n}_{k_{n}}}\Big($ $R^{(1)}_{S^{1}_{1...k_{1}}}\otimes...\otimes R^{(n)}_{S^{n}_{1...k_{n}}}\,\cdot$ (1)
$\mathds{1}_{S^{1}_{1...k_{1}-1},...,S^{n}_{1...k_{n}-1}}\otimes\tau_{S^{1}_{k_{1}},...,S^{n}_{k_{n}}}\Big)$
@@
has nonnegative expectation value on separable $\kappa$-partite states $\varrho_{S^{1}_{1...k_{1}-1},...,S^{n}_{1...k_{n}-1}}$, $\operatorname{tr}(\mathcal{W}_{\tau}\varrho)\geq 0$, where $\kappa=k_{1}+...+k_{n}-n$.
@@
This can be seen from the duality between the block-positive cone and the separable cone [21] as follows.
@@
Using the coordinate-free definition of the partial trace, for any state $\sigma^{(j)}_{k}$ and product state $\varrho^{(j)}=\varrho_{1}^{(j)}\otimes...\otimes\varrho_{k-1}^{(j)}$ we have $0\leq\operatorname{tr}(R^{(j)}\cdot\varrho^{(j)}\otimes\sigma^{(j)}_{k})=\operatorname{tr}(\theta^{(j)}\otimes\sigma^{(j)}_{k})$, where we define $\theta^{(j)}:=\operatorname{tr}_{1...k-1}(R^{(j)}\cdot\varrho_{1}^{(j)}\otimes...\otimes\varrho_{k-1}^{(j)}\otimes\mathds{1})$. Equivalently, $\theta^{(j)}$ is positive semidefinite for all $1\leq j\leq n$. Therefore, for any $n$-partite state or witness $\tau$, we have the inequality $0\leq\operatorname{tr}(\theta^{(1)}\otimes...\otimes\theta^{(n)}\cdot\tau)=\operatorname{tr}(\mathcal{W}_{\tau}\varrho)$. Here the equality is obtained by plugging in the definition (1) of $\mathcal{W}_{\tau}$, and we assume $\varrho:=\varrho^{(1)}\otimes...\otimes\varrho^{(n)}$ where each $\varrho^{(j)}$ is a product $(k_{j}-1)$-partite state. By convexity, $\operatorname{tr}(\mathcal{W}_{\tau}\varrho)\geq 0$ also holds for any separable $(k_{1}+...+k_{n}-n)$-partite state $\varrho$.
By the Choi-Jamiołkowski isomorphism, each state or witness opertor $R^{(j)}$ can be seen as the Choi operator of a map $\Psi_{j}(\varrho^{T}):=\operatorname{tr}_{1...k_{j}-1}(R^{(j)}\cdot\varrho_{1...k_{j}-1}\otimes\mathds{1}_{k_{j}})$ from systems $1...{k_{j}-1}$ to the system $k_{j}$}, whose output is positive semidefinite if the inputs are products of positive semidefinite operators. Further discussion focused on the case of trace polynomials is given in [22].
It is worth mentioning that this technique is not limited to using full-separability witnesses, in which case $\mathcal{W}_{\tau}$ is a full-separability witness as well. As an example beyond, we shall consider a bipartite state or witness $\tau_{A_{4}B_{4}}$ and two entanglement witnesses $W_{A_{1}A_{2}A_{3}A_{4}}$ and $V_{B_{1}B_{2}B_{3}B_{4}}$ detecting non-$2|1|1$-separability (i.e. have nonnegative expectation value on states where two parties are entangled and the rest are separable). In particular, let $W$ and $V$ have a nonnegative expectation value on four partite states of the form $\varrho_{1234}=\varrho_{123}\otimes\varrho_{4}$ and convex combinations thereof, where $\varrho_{123}=p_{1}\varrho_{1}\otimes\varrho_{23}+p_{2}\varrho_{2}\otimes\varrho_{13}+p_{3}\varrho_{12}\otimes\varrho_{3}$ with $p_{1},\,p_{2},\,p_{3}\geq 0$ and $p_{1}+p_{2}+p_{3}=1$ (namely $\varrho_{123}$ is biseparable). This gives rise to the $(\kappa=6)$-partite entanglement witness
$\mathcal{W}_{\tau}=\operatorname{tr}_{A_{4}B_{4}}\Big($ $W_{A_{1}A_{2}A_{3}A_{4}}\otimes V_{B_{1}B_{2}B_{3}B_{4}}\cdot$ (2)
$\quad\quad\quad\mathds{1}_{A_{1}A_{2}A_{3}B_{1}B_{2}B_{3}}\otimes\tau_{A_{4}B_{4}}\Big)\,.$
This witness has a nonnegative expectation value on states of the form $\varrho_{A_{1}A_{2}A_{3}B_{1}B_{2}B_{3}}=\sum_{i}q_{i}\varrho^{(i)}_{A_{1}A_{2}A_{3}}\otimes\varrho^{(i)}_{B_{1}B_{2}B_{3}}$ where $q_{i}\geq 0$, $\sum_{i}q_{i}=1$, and $\varrho^{(i)}_{A_{1}A_{2}A_{3}}$ and $\varrho^{(i)}_{B_{1}B_{2}B_{3}}$ are biseparable, and detects states out of this set if $\tau$ is chosen such that $\mathcal{W}_{\tau}$ has a negative eigenvalue.
@@
Figure 1: State-witness contraction yielding a four-partite witness from three- and bi- partite states and witnesses, with three-partite systems $S^{1}:=A$ and $S^{2}:=B$. First we take the tensor product of a witness $W$ and a positive operator $M$. Then the first (or last, equivalently) subsystems $A_{1}$ and $B_{1}$ are contracted (multiplied and partial traced) with a state or witness $\tau_{A_{1}B_{1}}$ to be chosen – see Eqs. (1) and (4). This allows to lift $W$ to a larger witness $\mathcal{W}_{\tau}$ with further detection capabilities (Observation 2 and Proposition 3), which moreover can be partially optimized over $\tau$ to detect specific four-partite states – see Eq. (6).
@@
Locally bound entanglement.    As a further case of interest, we will employ this method to reuse existing highly symmetric witnesses given in [23], which belong to a larger class of methods involving permutations in the Hilbert space [24, 25]. While these are desired due to their symmetries and implementation, their use in detecting states with local positive partial transpositions is a current challenge [26, 14, 23, 27, 28, 29, 30, 31]. In particular, consider the witness
@@
constructed from the so-called Young projectors $P_{2,1}=\eta_{d}\big(2(\operatorname{id})-(123)-(132)\big)/6$ and $P_{1^{3}}=\eta_{d}\big((\operatorname{id})-(12)-(13)-(23)+(123)+(132)\big)/6$, where $\eta_{d}(\pi)\mathinner{|v_{1}\rangle}\mathinner{|v_{2}\rangle}\mathinner{|v_{3}\rangle}=\mathinner{|v_{\pi^{-1}(1)}\rangle}\mathinner{|v_{\pi^{-1}(2)}\rangle}\mathinner{|v_{\pi^{-1}(3)}\rangle}\in{\mathds{C}^{d}}^{\otimes 3}$ represents the permutation $\pi\in S_{3}$ onto $(\mathds{C}^{d})^{\otimes 3}$} and $(\operatorname{id})$ is the identity permutation. Here permutations are written in cyclic notation and Young projectors $P_{\lambda}$ are labeled by partitions $\lambda\vdash n$ (see Appendix A for further details). This witness can be derived from immanant inequalities [32, 33] using the techniques in [23, 26, 20]. Linear optimization shows that it can detect three-partite states with zero, one or two positive partial transpositions, but not three (see Table II in [22]). Using Proposition 1 with an ancillary Bell state $\mathinner{|\phi^{+}\rangle}=\sum_{i=0}^{2}\mathinner{|ii\rangle}/\sqrt{3}$, it can be recycled to detect four-partite states whose local parties have positive partial transpositions, and thus cannot be distilled with local operations with classical communication if the four parties are spatially separated. This is a type of unlockable bound entanglement [34, 35, 36, 37, 38, 39], which we refer to as locally bound.
@@
can detect non-fully-separable four-partite states $\varrho_{A_{1}A_{2}B_{1}B_{2}}$ with $\varrho^{T_{i}}\geq 0$ for $i\in\{A_{1},A_{2},B_{1},B_{2}\}$. In particular, $\min\{\operatorname{tr}(\mathcal{W}_{\tau}):\varrho^{T_{i}}\geq 0\}\approx-0.000276$ .
Note that this can be achieved using the shared quantum resources provided by the Bell state in an ancillary system $A_{3}B_{3}$. Yet, these can be reduced significantly: using a less entangled state $\mathinner{|\phi_{s}\rangle}=\sqrt{s}\mathinner{|00\rangle}+\sqrt{1-s}\mathinner{|11\rangle}$ embedded into $\mathds{C}^{3}\otimes\mathds{C}^{3}$ with $s=0.9999$, a minimum value $\min_{\varrho^{T_{i}}\geq 0}\langle\mathcal{W}_{\tau}\rangle_{\varrho}\approx-6.051\cdot 10^{-5}$ is obtained using only $s\log_{2}(s)-(1-s)\log_{2}(1-s)\approx 0.0015$ ebits of entanglement.
Multipartite bound entanglement.   Here we consider the detection of multipartite entangled states that cannot be distilled to pure entanglement across any of the bipartitions. This is the case of multipartite entangled states whose partial transpositions across all bipartitions are positive semidefinite (MPPT). Our goal is to lift non-decomposable multipartite witnesses into larger ones, also non-decomposable. Namely, to prove the following:
Proposition 3.
Using SWC, it is possible to use an entanglement witness detecting $k$-partite MPPT states of certain size, to create a new witness detecting $k^{\prime}>k$-partite MPPT states.
The proof shall be shown with an analytical example. Consider a $k$-partite witness written in the computational basis, $W=\sum_{i_{1},j_{i},...,i_{k},j_{k}=0}^{d-1}w_{i_{1}j_{1}...i_{k}j_{k}}\mathinner{|i_{1}\rangle}\mathinner{\langle j_{1}|}\otimes...\otimes\mathinner{|i_{k}\rangle}\mathinner{\langle j_{k}|}$; and the unnormalized $l$-partite GHZ state $\mathinner{|{\rm{GHZ}}_{l,d}\rangle}=\sum_{t=0}^{d-1}\mathinner{|t\rangle}^{\otimes l}$. We shall use SWC with $M=|{\rm{GHZ}}_{l,d}\rangle\langle{\rm{GHZ}}_{l,d}|$ to obtain a $\kappa$-partite witness $\mathcal{W}=\operatorname{tr}_{k,k+l}\big(W_{1...k}\otimes M_{k+1...k+l}\cdot\mathds{1}\otimes\tau_{k,k+l}\big)$, where $\kappa=k+l-2$. By employing $\tau$ to be an unnormalized maximally entangled state $d|\phi^{+}\rangle\langle\phi^{+}|$ or the SWAP operator $\eta_{d}(12)=d|\phi^{+}\rangle\langle\phi^{+}|^{T_{2}}$, one obtains respectively the witness
$\mathcal{W}_{\phi^{+}}=\sum_{\vec{i},\vec{j}\in\mathbf{Z}_{d}^{k}}w_{\vec{i},\vec{j}}\mathinner{|i_{1}\rangle}\mathinner{\langle j_{1}|}\otimes...\otimes\mathinner{|i_{k}\rangle}\mathinner{\langle j_{k}|}^{\otimes l-1}$ (5)
with $\vec{i}=i_{1},...,i_{k}$ and $\vec{j}=j_{1},...,j_{k}$, or its partial transposition $\mathcal{W}_{(12)}=\mathcal{W}_{\phi^{+}}^{T_{k,...,k+l-2}}$. If the $k$-partite witness used is the three-qubit witness $W_{I}$ dectecting locally PPT states given in [40], then we can find a new entanglement witness $\mathcal{W}_{\phi^{+}}$ detecting MPPT states shared among four and five qubits. This is detailed in Appendix B, together with the obtention of less sparse witnesses for MPPT states through different choices of the operator $M$.
Witnesses tailored to states.   While in Proposition 1 one can choose $\tau$ to be any block-positive operator (state or witness), the optimization becomes accessible with semidefinite programming if $\tau$ is restricted to be a state or a decomposable witness [41] (or indecomposable, within the first levels of symmetric extension [12, 42]). Consider a fixed set of $n$ $k_{j}$-partite witnesses (or states) $R^{(i)}$ and a target $(\kappa=k_{1}+...+k_{n}-n)$-partite state $\varrho$ in local dimension $d$ to be detected. Let us choose $\tau$ to be optimized over all possible states and decomposable witnesses [9],
@@
Here $\mathcal{S}$ denotes any possible subset of the last subsystems $\{k_{1},...,k_{n}\}$ and labels the operators $X_{\mathcal{S}}$ acting on $\mathds{C}^{d_{k_{1}}}\otimes\cdots\otimes\mathds{C}^{d_{k_{n}}}$, and thus the second line imposes that $\tau$ is a decomposable $n$-partite entanglement witness (or a state, if optimality is achieved for $X_{\mathcal{S}}=0$). Notice that the variable $\tau$ is of size $d^{n}$ [43], even though the target state to be detected is described by a density matrix of size $d^{(k-1)n}$ (assuming $k_{1}=...=k_{n}:=k$ for simplicity).
@@
$\mathcal{W}_{\tau^{*}}=\operatorname{tr}_{A_{2}B_{2}}\Big(\eta_{d}(12)_{A_{1}A_{2}}\otimes\phi^{+}_{B_{1}B_{2}}\cdot\varrho_{A_{1}B_{1}}\otimes\tau^{*}_{A_{2}B_{2}}\Big)$ (7)
with $\tau^{*}=\text{argmin}_{\tau}\big\{\operatorname{tr}(\mathcal{W}_{\tau}\varrho):\tau\geq 0,\operatorname{tr}\tau=1\big\}$ detects any state $\varrho$ with nonpositive partial transpose. Similarly, standard witness tailoring techniques using diagonalization [9] are recovered with this example: the duality between positive maps and witnesses [44] is used to detect any state detected by a positive map $\Lambda_{A}\otimes\operatorname{id}_{B}(\varrho_{AB})\not\geq 0$, by optimizing over $\tau^{*}$ and replacing $\phi^{+}$ in Eq. (7) for $\Lambda^{\dagger}\otimes\operatorname{id}(\eta_{d}(12))$ where $\Lambda^{\dagger}$ satisfies $\operatorname{tr}(\Lambda(X)Y)=\operatorname{tr}(X\Lambda^{\dagger}(Y))$ for all operators $X$ and $Y$.
Multiple copies.   Detecting entanglement in copies of a state can be very advantageous with respect to linear methods [19, 18, 17]. However, nonlinearity makes it hard to tailor such techniques to desired states. Here we show how Eq. (6) allows to partially solve this problem with simple semidefinite program whose size is only the size of the state. As a case study, we shall significantly improve our previous trace-polynomial multicopy constructions given in [20]: notice that the latter are obtained by setting in Proposition 1 the input state to be composed of k1k-1 copies of a $\kappa$-partite target state $\varrho^{\otimes k-1}$ with $\kappa=2(k-1)$ and the naive choice $\tau=\varrho$. For example, consider a Bell state affected by white noise,
@@
which is often used as a worse-case scenario in noisy setups [45]. In [20] we showed that this state can be detected for a range of values of $p$ with the nonlinear witness $\mathcal{W}=(\mathds{1}-6P_{1^{3}})_{A_{123}}\otimes(P_{1^{3}})_{B_{123}}$ acting on three copies of $\varrho$, through $0<\operatorname{tr}(\mathcal{W}\varrho^{\otimes 3})=\operatorname{tr}(\mathcal{W}_{\varrho}\varrho^{\otimes 2})$ where we denote $\mathcal{W}_{\varrho}:=\operatorname{tr}_{A_{3}B_{3}}(\mathcal{W}\cdot\mathds{1}\otimes\varrho_{A_{3}B_{3}})$ for easier comparison to our present work. From Proposition 1 it becomes clear that $\mathcal{W}$ can be better tuned to $\varrho$ using the freedom $\tau\neq\varrho$ given by Eq. (6), in this case leading to (Fig. 1)
$).\mathcal{W}_{\tau}:=\operatorname{tr}_{A_{3}B_{3}}\Big(\mathcal{W}\cdot(\mathds{1}_{A_{1}A_{2}B_{1}B_{2}}\otimes\phi^{+}_{A_{3}B_{3}}$$\Big)\,$. (9)
@@
Figure 2: Tailoring multicopy witnesses to states. Linear witnesses detect regions of entangled states separated by hyperplanes (magenta area) through $\langle W\rangle_{\varrho}=\operatorname{tr}(W\varrho)<0$. Multicopy witnesses $\mathcal{W}$ detect regions of entangled states through $\langle\mathcal{W}\rangle_{\varrho^{\otimes k}}<0$ [18, 19], separated by curved hypersurfaces (purple and blue areas). In particular, $\mathcal{W}$ can be constructed by tensoring linear witnesses [20]. By replacing one of the copies by a variable state or witness $\tau$, the state-witness contraction technique (SWC) allows to systematically detect a larger set of states (green area –see Table 1 for an example) and approach further the set of separable states (SEP).
@@
Observation 4.
Let $\mathcal{W}_{\tau}$ in Eq. (6) be constructed from trace polynomial operators $R^{(j)}$ obtained by immanant inequalities according to [23]. Then, the minimization of $\operatorname{tr}(\mathcal{W}_{\tau}\varrho^{\otimes k-1})$ over $\tau$ is invariant under local unitaries acting on $\varrho$.
@@
Local unitaries act as $U^{\otimes k-1}$ onto the first k1k-1 copies of each local party. A minimum obtained for $\varrho_{AB}$ by $\tau_{AB}$ is also obtained for $U\otimes V\varrho_{AB}U^{\dagger}\otimes V^{\dagger}$ by $U\otimes V\tau_{AB}U^{\dagger}\otimes V^{\dagger}$ (extension to the multipartite case is straightforward). Thus local unitaries act in practice as $U^{\otimes k}$ on each party. The rest of the argument is the same as in [20], namely $U^{\otimes k}$ commutes with the representation $\eta_{d}(\pi)$ of permutations $\pi\in S_{k}$ due to the Schur-Weyl duality. ∎